OpenStudy (anonymous):

solve for x. (1/a)+(1/x)=(1/b)

hartnn (hartnn):

subtract 1/a from both sides, what u get?

OpenStudy (anonymous):

(1/x) = (1/b)-(1/a)

OpenStudy (anonymous):

1/x = 1/b -1/a or 1/x =(a-b)/(ab) as LCM of denom is ab of rhs then x =(ab)/(a-b)

hartnn (hartnn):

@matricked stop giving entire solutions

hartnn (hartnn):

can u simplify 1/b-1/a ?

OpenStudy (anonymous):

no....?

hartnn (hartnn):

use this \(\huge\frac{p}{q}-\frac{r}{s}=\frac{ps-rq}{qs}\)

OpenStudy (jiteshmeghwal9):

ok ! hint :- \[\frac{(1 \times a)-(1 \times b)}{ba}\]

OpenStudy (jiteshmeghwal9):

now what do u get @spasta106 ??

OpenStudy (anonymous):

a-b)/ba ?

OpenStudy (jiteshmeghwal9):

ok ! so \[\frac{1}{x}=\frac{a-b}{ba}\]now cross multiply what do u gt @spasta106 ??

hartnn (hartnn):

or just flip the fraction....

OpenStudy (anonymous):

x(a-b)=ba

OpenStudy (anonymous):

so x = (ba)/(a-b)

hartnn (hartnn):

yup

OpenStudy (jiteshmeghwal9):

yup ! @hartnn is also correct :)

OpenStudy (jiteshmeghwal9):

gt it @spasta106 ?

OpenStudy (anonymous):

\[\frac{ 1 }{ x } = \frac{ 1}{ b } - \frac{ 1 }{ a }\] \[\frac{ 1 }{ x } = \frac{ \left( a-b \right) }{ ab }\] \[x = \frac{ ab }{ \left( a-b \right) }\]